OpenStudy (anonymous):

-2(v-7)-6v=8(4-v)-3 help?

OpenStudy (anonymous):

-2(v-7)-6v=8(4-v)-3 Divide both sides by -2 (v-7)-6v=(8(4-v)-3)/2 Simplify (v-7)-6v=4(4-v)-3/2 -5v-7=8-4v-3/2 Sum both sides 4v -v-7=8-3/2 sum both sides 7 -v=27/2 Multiply by minus 1 both sides v=-27/2 That is your answer, you cant prove by put it into the equation. And is false, is not a equality, is false.